# Multilevel neural simulation-based inference

### v2

@@ 4 Theory
Suppose the assumptions of Theorem˜1 hold. Then, the values of $n_{0},\ldots,n_{L}$ which minimise the upper bound on $\text{Var}[{\ell}_{\text{MLMC}}(\phi)]$ in ˜5 given a fixed computational budget; i.e. for $\text{Cost}({\ell}_{\text{MLMC}}(\phi);n_{0},\ldots,n_{L})\leq C_{\text{budget}}$ for some $C_{\text{budget}}>0$ are given by
$n^{\star}_{0}\propto\frac{C_{\text{budget}}}{\sqrt{C_{0}}}\sqrt{\left\|G^{0}\right\|_{W^{1,4}(\pi\times\mathbb{U}^{m})}^{4}+1},\quad n^{\star}_{l}\propto\frac{C_{\text{budget}}}{\sqrt{C_{l}+C_{l-1}}}\sqrt{\left\|G^{l}-G^{l-1}\right\|^{2}_{W^{1,4}(\pi\times\mathbb{U}^{m})}+1}.$
See Section˜A.2 for the proof. This result provides useful intuition on how to select the number of sample per level. For instance, consider the case $L=1$, i.e., two levels. If $G^{0}$ is known to be a good approximation of $G^{1}$, it makes sense to allocate a large budget to generating low-fidelity simulations while only a small number of high-fidelity simulations would be sufficient. On the other hand, if $G^{0}$ and $G^{1}$ differ substantially, allocating a larger budget to high-fidelity simulations makes sense, despite the higher cost, as it can help accurately capture this difference. Although Theorem˜2 provides intuition, it may be hard to obtain the optimal $n_{0},\dots,n_{L}$ exactly since it requires computing quantities which are often unknown. For example, computing Sobolev norms can be challenging, and the results implicitly depend on $\phi$ through the constants $K_{0}(\phi),\ldots,K_{L}(\phi)$ of Theorem˜1 (which are unlikely to be tight). This is a common limitation of MLMC theory; see (Giles, 2015, Sec. 2-3).
@@ 5.2 Ornstein–Uhlenbeck process: a popular financial model
Our next experiment involves the Ornstein-Uhlenbeck (OU) process—a stochastic differential equation model commonly used in financial analysis (Minenna, 2003), which in our case outputs a $100$-dimensional Markovian time-series and has three parameters. The process is known to converge to a stationary Gaussian distribution, which we use as the low-fidelity simulator, see Section˜B.3. The example is reproduced from Krouglova et al. (2025), who used it for benchmarking their transfer learning approach to NPE (TL-NPE). TL-NPE first trains an NPE network on low-fidelity simulations, and then uses a small set of high-fidelity data to refine the network parameters until a stopping criterion is met. We implement TL-NPE with $n_{0}=1100$ and $n_{1}=100$, using 10% of data as validation set for their stopping criterion, and compare against ML-NPE with $n_{0}=1000$ and $n_{1}=100$. These values are picked so as to keep the simulation budget same for both the methods.
We compare ML-NPE with TL-NPE for different choices of its hyperparameter in Figure˜3. TL-NPE’s hyperparameter is the number of epochs until the stopping criterion is activated, termed patience, which we let take values in $\{1,20,10^{2},10^{3}\}$. We use two metrics for evaluation: the NLPD of the true $\theta$ under the estimated posterior, and the KLD between the posterior obtained using either TL-NPE or ML-NPE and a reference NPE posterior trained using $n=10,000$ samples. We observe that TL-NPE’s performance is quite sensitive to the choice of the patience value, and the authors of that work do not provide guidance on how to select this in practice (their code fixes this to $20$ with no clear justification). When the number of high-fidelity samples are low ($n_{1}=10$), its performance degrades as patience increases, as a higher patience means TL-NPE is overfitting to the few high-fidelity samples, losing the information gained from training on the low-fidelity data. In contrast, ML-NPE achieves better performance especially when $n_{1}=10$ and does not have any hyperparameter to tune, making it easier to use in practice. We additionally investigate a challenging scenario where the high- and low-fidelity simulators have differing dimensionalities of $\theta$, in which our method exhibits some limitations, possibly due to instability during optimisation; see Section˜B.3.1 for details.
@@
Figure 3: NLPD and KLD for ML-NPE (ours) and TL-NPE (Krouglova et al., 2025) using different patience values (hyperparameter of TL-NPE) over 2020 runs. ML-NPE performs better in the limited high-fidelity data scenario ($n_{1}=10$) (a), whilst being competitive when $n_{1}=100$ (b). TL-NPE shows sensitivity towards its hyperparameter, while ML-NPE does not require any hyperparameter.
@@ 5.3 Toggle-switch model: a Systems Biology example
We now consider the toggle-switch model (Bonassi et al., 2011; Bonassi and West, 2015). This model describes the interaction between two genes over time, and has seven parameters and a scalar observation at the end of a time interval. Simulators typically use time-discretisation, and the total number of time-steps $T$ acts as a fidelity parameter: running the model with large $T$ incurs larger computational cost but leads to accurate simulations, while smaller $T$ leads to cheap but inaccurate samples. Thus, this model illustrates a setting with more than two fidelity levels (by taking $T$ to be more than two values)—a situation straightforward to handle with our multilevel approach but not with TL-NPE (Krouglova et al., 2025). We take $T_{0}=50$, $T_{1}=80$, and $T_{2}=300$ to be the number of steps for the three fidelity levels with $n_{0}=10^{4}$, $n_{1}=500$, and $n_{2}=100$. Guided by the intuition from Theorem˜2, we allocate a large budget to estimating the difference between the low- and the medium-fidelity simulators, leveraging prior knowledge that this difference is substantial. See Section˜B.4.1 for results under alternative budget allocations. Note that in this case each fidelity level has a different base measure of dimension $2T+1$; however, there is still sufficient seed-matching (see Section˜B.4 for details), and therefore variance reduction, thanks to MLMC. We compare our ML-NLE method with the standard NLE trained on samples from either the low- ($n=12,060$), the medium- ($n=7537$), or the high-fidelity simulator ($n=2010$). The number of training data $n$ in each case is selected so as to match the total computational cost of simulation between ML-NLE and NLE, see Section˜B.4. Figure˜4 reports the maximum mean discrepancy (MMD) (Gretton et al., 2012) between $500$ samples from the learned conditional densities and $500$ samples from the high-fidelity simulator across for $5000$ different parameter values. We observe that ML-NLE performs better than all the NLE baselines for the same computational cost.
@@ 5.4 Cosmological Simulations
We now consider a cosmological simulator using the CAMELS suite (Villaescusa-Navarro et al., 2021, 2023)—one of the most computationally intensive cosmological simulations to date—which comprises both low and high fidelity data. These are real state-of-the-art simulations being used with real-world observational data. Our task is to infer a standard cosmological target parameter using a 39-dimensional power spectra of cosmology data, see Section˜B.5 for an example. The low-fidelity simulations are gravity-only N-body simulations, whose physical behaviour is controlled only by the parameter and some Gaussian fluctuations of the initial conditions. The high-fidelity hydrodynamic simulations have additional physics, controlled by an additional five parameters. We include these additional parameters as part of the $\mathcal{U}$ space, constituting a case of partial common random numbers between the low- and high-fidelity simulators, similar to Section˜5.3.
@@
Here, the high-fidelity simulations can be orders of magnitude ($\times 100$) slower to generate than the low-fidelity ones, making this a representative problem for our method. Assuming we only have access to $n=n_{1}=20$ high-fidelity simulations, we wish to ascertain the improvement in inference accuracy by including 1000 low-fidelity simulations (i.e. $n_{0}=980$) using our ML-NPE method. To that end, we measure the NLPD and empirical coverage of the estimated posteriors for 980 test data. The result in Figure˜5 shows that ML-NPE performs better than standard NPE for both the metrics. Standard NPE tends to produce overconfident posteriors, while ML-NPE yields calibrated or underconfident posteriors for most confidence levels. Thus, including low-fidelity samples using our method leads to better inference outcomes. Before concluding, we note that several papers demonstrating the potential of multi-fidelity SBI methods in cosmology appeared around the same time as our paper (Saoulis et al., 2025; Thiele et al., 2025). These more in-depth studies clearly highlight the potential for impact of advanced multi-fidelity SBI methods.
@@ 6 Conclusion
This paper demonstrated how to reduce the cost of SBI using MLMC, but could more broadly be seen as a way to perform multilevel training of conditional density estimators in scenarios where data from different sources with different accuracy levels needs to be combined. Our method achieves this without any additional fidelity hyperparameter to tune and can handle more than two fidelity levels. It is also particularly appealing since it is complementary to other compute-efficient SBI methods. For example, Tatsuoka et al. (2025) recently proposed to train an NPE network on low-fidelity data, and to then use the resulting posterior approximation to guide sampling from the high-fidelity simulator. This approach could easily be combined with our method.
@@ Acknowledgments
The authors are grateful to Tim Sullivan and David Warne for helpful discussions, and to the authors of Krouglova et al. (2025) for sharing their source code. YH and AB were supported by the Research Council of Finland grant no. 362534. FXB was supported by the EPSRC grant [EP/Y022300/1]. NJ was supported by the ERC-selected UKRI Frontier Research Grant EP/Y03015X/1 and by the Simons Collaboration on Learning the Universe.
@@ A.1 Proof of Theorem˜1
Note that here, and throughout the rest of this proof, we will simplify the notation. We will drop the subscript for variances and expectations, and these should all be understood as being $u_{1:m}\sim\mathbb{U}^{m}$ and $\theta\sim\pi$. We will also use the variable $z$ to denote a vector containing $u_{1:m}$ and $\theta$, so that $z\in\mathbb{R}^{md_{\mathcal{U}}+d_{\Theta}}$ and $z_{i}$ represents a vector containing $u_{1:m,i}$ and $\theta_{i}$. Finally, we will write $L^{4}$ and $W^{1,4}$ to denote the weighted norms $L^{4}(\pi\times\mathbb{U}^{m})$ and $W^{1,4}(\pi\times\mathbb{U}^{m})$.
The proof will be structured as follows. We will express the variance of each term using the variance of individual samples. We will then use a Poincaré-type inequality to bound the variance of the objective in terms of the expected squared norm of its gradient, then we will upper bound the norm using only terms which depend on constants and terms expressing how well $G^{l-1}$ approximates $G^{l}$. Note that we will first prove the result for NLE, and will comment on the NPE case at the bottom of the proof.
Since each term in the MLMC expansion are MC estimators (and therefore based on independent samples), we can express the variances of each term as follows:
@@
for $l\in\{1,\ldots,L\}$. Here, we emphasise again that the expectation is over $z$, which encompasses both $\theta$ and $u_{1:m}$. This result requires the joint distribution of the prior $\pi$ and $m$ times the base measure $\mathbb{U}$ to be log-concave, which holds since the product of log-concave densities is also log-concave (since the sum of convex functions is convex) and we have assumed that the prior and base measures are independent and separately log-concave through Assumption (A1).
To simplify notation, we now introduce the vector-valued function $g^{l}(z)=g^{l}(\theta,u_{1:m})=(G^{l}_{\theta}(u_{1})^{\top},\ldots,G^{l}_{\theta}(u_{m})^{\top},\theta)^{\top}$ so that $q_{\phi}^{\text{NLE}}(G^{l}_{\theta}(u)\,|\,\theta)=q_{\phi}^{\text{NLE}}(g^{l}(\theta,u))=q_{\phi}^{\text{NLE}}(g^{l}(z))$. We will now derive our first bound, which looks at the first term in the MLMC expansion. To do so, we simplify ˜8 as follows
$\mathbb{E}\left[\left\|\nabla_{z}f_{\phi}^{0}(z)\right\|^{2}_{2}\right]$ $=\mathbb{E}\left[\left\|\nabla_{z}\log q^{\text{NLE}}_{\phi}\left(g^{0}(z)\right)\right\|^{2}_{2}\right]$
$=\mathbb{E}\left[\left\|\nabla_{z}g^{0}(z)\nabla\log q^{\text{NLE}}_{\phi}\left(g^{0}(z)\right)\right\|^{2}_{2}\right]$ (10)
$\leq\mathbb{E}\left[\left\|\nabla_{z}g^{0}(z)\right\|^{2}_{2}\left\|\nabla\log q^{\text{NLE}}_{\phi}\left(g^{0}(z)\right)\right\|^{2}_{2}\right]$ (11)
$\leq\mathbb{E}\left[\left\|\nabla_{z}g^{0}(z)\right\|^{4}_{2}\right]^{\frac{1}{2}}\mathbb{E}\left[\left\|\nabla\log q^{\text{NLE}}_{\phi}\left(g^{0}(z)\right)\right\|^{4}_{2}\right]^{\frac{1}{2}}$ (12)
Here, we have that ˜10 follows due to the chain rule, ˜11 follows due to the Cauchy-Schwarz inequality of the 2-norm, ˜12 follows due to the Cauchy-Schwarz inequality of expectations.
@@
$\mathbb{E}\left[\left\|\nabla_{z}g^{0}(z)\right\|^{4}_{2}\right]$ $=2\mathbb{E}\left[\|\nabla_{z}G^{0}(z)\|^{4}_{2}\right]+2\mathbb{E}\left[\|\nabla_{z}\theta\|_{2}^{4}\right]$ (13)
$\leq 2\mathbb{E}\left[\left\|\nabla_{z}G^{0}(z)\right\|^{4}_{F}\right]+2\mathbb{E}\left[\|\nabla_{z}\theta\|_{F}^{4}\right]$ (14)
$=2\mathbb{E}\left[\left(\sum_{|\alpha|=1}\left(\sum_{j=1}^{d}\left(D^{\alpha}(G^{0}(z))_{j}\right)^{2}\right)\right)^{2}\right]+2d_{\Theta}^{2}$ (15)
$\leq 2(d_{\Theta}+md_{\mathcal{U}})\sum_{|\alpha|=1}\mathbb{E}\left[\left(\sum_{j=1}^{d}\left(D^{\alpha}(G^{0}(z))_{j}\right)^{2}\right)^{2}\right]+2d_{\Theta}^{2}$ (16)
$\leq 2(d_{\Theta}+md_{\mathcal{U}})\sum_{|\alpha|=1}\mathbb{E}\left[\|D^{\alpha}G^{0}(z)\|_{2}^{4}\right]+2d_{\Theta}^{2}$ (17)
$\leq 2(d_{\Theta}+md_{\mathcal{U}})\sum_{|\alpha|=1}\left\|D^{\alpha}G^{0}\right\|^{4}_{L^{4}}+2d_{\Theta}^{2}$ (18)
$\leq 2(d_{\Theta}+md_{\mathcal{U}})\left\|G^{0}\right\|^{4}_{W^{1,4}}+2d_{\Theta}^{2},$ (19)
@@
where ˜13 follows from the fact that $(a+b)^{2}\leq 2a^{2}+2b^{2}$, ˜14 follows from the fact that the Frobenius norm $\|\cdot\|_{F}$ is an upper-bound on the matrix 2-norm, ˜15 follows from the definition of the Frobenius norm and from the fact that $\mathbb{E}[\|\nabla_{z}\theta\|_{F}^{4}]=d_{\Theta}^{2}$, ˜16 follows from the fact that $(\sum_{i=1}^{n}a_{i})^{2}\leq n\sum_{i=1}^{n}a_{i}^{2}$, and ˜17, ˜18 and ˜20 follow from the definitions of the 2-norm, the $L^{4}$-norm and the $W^{1,4}$-norm respectively.
We now move on to bounding the second term in ˜12. Since we assumed in Assumption (A3) that $\nabla\log q^{\text{NLE}}_{\phi}$ satisfies a linear growth condition, we must have that
$\mathbb{E}\left[\left\|\nabla\log q^{\text{NLE}}_{\phi}\left(g^{0}(z)\right)\right\|^{4}_{2}\right]$ $=\mathbb{E}\left[\left\|\nabla\log q^{\text{NLE}}_{\phi}\left(G_{\theta}^{0}(u)\,|\,\theta\right)\right\|^{4}_{2}\right]$ (21)
$\leq K^{4}_{\text{grow}}(\phi)\mathbb{E}\left[\left(\left\|G^{0}(z)\right\|_{2}+\|\theta\|_{2}+1\right)^{4}\right]$ (22)
$\leq 27K^{4}_{\text{grow}}(\phi)\left(\mathbb{E}\left[\left\|G^{0}(z)\right\|_{2}^{4}\right]+\mathbb{E}\left[\|\theta\|_{2}^{4}\right]+1\right)$ (23)
$\leq 27K^{4}_{\text{grow}}(\phi)\left(\left\|G^{0}\right\|_{W^{1,4}}^{4}+\mathbb{E}\left[\|\theta\|_{2}^{4}\right]+1\right)$
$\leq K_{\text{score}}(\phi)\left(\left\|G^{0}\right\|_{W^{1,4}}^{4}+1\right)$ (24)
Here, ˜22 uses the growth condition, ˜23 holds due since $(a+b+c)^{4}\leq 27(a^{4}+b^{4}+c^{4})$, and ˜24 uses the definition of Sobolev norm and the fact that $\mathbb{E}\left[\|\theta\|_{2}^{4}\right]$ is upper bounded by a constant since the expectation is against $\pi$, which a log-concave distribution and hence all its moments are finite (see Section 5.1. of Saumard and Wellner [2014]). Combining Section˜A.1, ˜8, ˜12, ˜20, and ˜24 therefore gives:

### v4

@@ 4 Theory
Suppose the assumptions of Theorem˜1 hold. Then, the values of $n_{0},\ldots,n_{L}$ which minimise the upper bound on $\text{Var}[{\ell}_{\text{MLMC}}(\phi)]$ in ˜5 given a fixed computational budget; i.e. for $\text{Cost}({\ell}_{\text{MLMC}}(\phi);n_{0},\ldots,n_{L})\leq C_{\text{budget}}$ for some $C_{\text{budget}}>0$, are given by
$n^{\star}_{0}\propto\frac{C_{\text{budget}}}{\sqrt{C_{0}}}\sqrt{\left\|G^{0}\right\|_{W^{1,4}(\pi\times\mathbb{U})}^{4}+1},\quad n^{\star}_{l}\propto\frac{C_{\text{budget}}}{\sqrt{C_{l}+C_{l-1}}}\left\|G^{l}-G^{l-1}\right\|_{W^{1,4}(\pi\times\mathbb{U})}.$
See Section˜A.2 for the proof. This result provides useful intuition on how to select the number of samples per level. For instance, consider the case $L=1$, i.e., two levels. If $G^{0}$ is known to be a good approximation of $G^{1}$, it makes sense to allocate a large budget to generating low-fidelity simulations while only a small number of high-fidelity simulations would be sufficient. On the other hand, if $G^{0}$ and $G^{1}$ differ substantially, allocating a larger budget to high-fidelity simulations makes sense, despite the higher cost, as it can help accurately capture this difference. Although Theorem˜2 provides intuition, it may be hard to obtain the optimal $n_{0},\dots,n_{L}$ exactly since it requires computing quantities which are often unknown. For example, computing Sobolev norms can be challenging, and the results implicitly depend on $\phi$ through the constants $K_{0}(\phi),\ldots,K_{L}(\phi)$ of Theorem˜1 (which are unlikely to be tight). This is a common limitation of MLMC theory; see (Giles, 2015, Sec. 2-3).
Before concluding this section, we note that our theory focussed on the variance of $\ell_{\text{MLMC}}(\phi)$, but another important quantity for gradient-based optimisation will be the variance of the gradient $\nabla_{\phi}\ell_{\text{MLMC}}(\phi)$ of this objective. It turns out that similar results are straightforward to prove for this quantity under very minor modifications of the assumptions, see Section˜A.3.
@@ 5.2 Ornstein–Uhlenbeck process: a popular financial model
Our next experiment involves the Ornstein-Uhlenbeck (OU) process—a stochastic differential equation model commonly used in financial analysis (Minenna, 2003), which in our case outputs a $100$-dimensional Markovian time-series and has three parameters. The process is known to converge to a stationary Gaussian distribution, which we use as the low-fidelity simulator, see Section˜B.3. The example is reproduced from Krouglova et al. (2025), who used it for benchmarking their transfer learning approach to NPE (TL-NPE). TL-NPE first trains an NPE network on low-fidelity simulations, and then uses a small set of high-fidelity data to refine the network parameters until a stopping criterion is met. We implement TL-NPE with $n_{0}=1100$ and $n_{1}=\{10,100\}$, and compare against ML-NPE with $n_{0}=1000$ and $n_{1}=\{10,100\}$. These values are picked to keep the simulation budget the same for both methods. TL-NPE uses the default early stopping criterion from the sbi library (Tejero-Cantero et al., 2020), which terminates training if the validation loss increases for $20$ epochs. Once the criterion is met, the network parameters achieving the lowest validation loss are selected. This criterion is used for both the low- and the high-fidelity training stages. We use 20% of the data as validation set for early stopping.
We run both methods $20$ times and compute the NLPD of the ground-truth parameter $\theta$ under the estimated posterior across $500$ test points. Additionally, we report the KLD between the posterior obtained using TL-NPE or ML-NPE and a reference NPE posterior trained with n=10,000n=10{,}000 high-fidelity simulations. Results are aggregated across the 20 runs and all test points, and reported in Figure˜8. We observe that both the methods achieve comparable performance: for $n_{1}=10$, ML-NPE performs slightly better on both metrics, whereas for $n_{1}=100$, TL-NPE slightly outperforms ML-NPE on both.
We additionally investigate a challenging scenario where the high- and low-fidelity simulators have differing dimensionalities of $\theta$, in which our method exhibits some limitations, possibly due to instability during optimisation; see Section˜B.3.1 for details.
@@
Figure 3: Performance of ML-NPE and TL-NPE measured by NLPD ($\downarrow$) and KL divergence ($\downarrow$) with different number of high-fidelity samples $n_{1}$ on the OU process. (a) When $n_{1}=10$, ML-NPE outperforms on both metric. (b) When $n_{1}=100$, TL-NPE outperforms on both metric. Note that for visualisation purpose, we removed outliers, see Section˜B.3 for the result with all the data.
@@ 5.3 Toggle-switch model: a Systems Biology example
We now consider the toggle-switch model (Bonassi et al., 2011; Bonassi and West, 2015). This model describes the interaction between two genes over time, and has seven parameters and a scalar observation at the end of a time interval. Simulators typically use time-discretisation, and the total number of time-steps $T$ acts as a fidelity parameter: running the model with large $T$ incurs larger computational cost but leads to accurate simulations, while smaller $T$ leads to cheap but inaccurate samples. Thus, this model illustrates a setting with more than two fidelity levels (by taking $T$ to be more than two values). We take $T_{0}=50$, $T_{1}=80$, and $T_{2}=300$ to be the number of steps for the three fidelity levels with $n_{0}=10^{4}$, $n_{1}=500$, and $n_{2}=100$. Guided by the intuition from Theorem˜2, we allocate a large budget to estimating the difference between the low- and the medium-fidelity simulators, leveraging prior knowledge that this difference is substantial. See Section˜B.4.1 for results under alternative budget allocations. Note that in this case each fidelity level has a different base measure of dimension $2T+1$; however, there is still sufficient seed-matching (see Section˜B.4 for details), and therefore variance reduction, thanks to MLMC. We compare our ML-NLE method with the standard NLE trained on samples from either the low- ($n=12,060$), the medium- ($n=7537$), or the high-fidelity simulator ($n=2010$). The number of training data $n$ in each case is selected so as to match the total computational cost of simulation between ML-NLE and NLE, see Section˜B.4. Figure˜4 reports the maximum mean discrepancy (MMD) (Gretton et al., 2012) between $500$ samples from the learned conditional densities and $500$ samples from the high-fidelity simulator across for $5000$ different parameter values. We observe that ML-NLE performs better than all the NLE baselines for the same computational cost.
@@ 5.4 Cosmological Simulations
We now consider a cosmological simulator using the CAMELS suite (Villaescusa-Navarro and others, 2021, 2023)—one of the most computationally intensive cosmological simulations to date—which comprises both low- and high-fidelity data. These are state-of-the-art simulations being used with real-world data. Developing surrogate, multilevel techniques for cosmology has become a key research focus (Chartier et al., 2021; Chartier and Wandelt, 2022). Our task is to infer a standard cosmological target parameter using a 39-dimensional power spectra of cosmology data, see Section˜B.5 for an example. The low-fidelity simulations are gravity-only N-body simulations, whose physical behaviour is controlled only by the parameter and some Gaussian fluctuations of the initial conditions. The high-fidelity hydrodynamic simulations have additional physics, controlled by an additional five parameters. We include these additional parameters as part of the $\mathcal{U}$ space, constituting a case of partial common random numbers between the low- and high-fidelity simulators, similar to Section˜5.3.
@@
Here, the high-fidelity simulations can be orders of magnitude ($\times 100$) slower to generate than the low-fidelity ones, making this a representative problem for our method. Assuming we only have access to $n=n_{1}=20$ high-fidelity simulations, we wish to ascertain the improvement in inference accuracy by including 1000 low-fidelity simulations (i.e. $n_{0}=980$) using our ML-NPE method. To that end, we measure the NLPD and empirical coverage of the estimated posteriors for 980 test data. The result in Figure˜5 shows that ML-NPE performs better than standard NPE for both the metrics. Standard NPE tends to produce overconfident posteriors, while ML-NPE yields calibrated or underconfident posteriors for most confidence levels. Thus, including low-fidelity samples using our method leads to better inference outcomes. Before concluding, we note that some recent papers demonstrating the potential of multi-fidelity SBI methods in cosmology appeared around the same time as our paper (Saoulis et al., 2025; Thiele et al., 2025). These more in-depth studies clearly highlight the potential for impact of advanced multi-fidelity SBI methods.
@@ 6 Conclusion
This paper demonstrated how to reduce the cost of SBI using MLMC, but could more broadly be seen as a way to perform multilevel training of conditional density estimators in scenarios where data from different sources with different accuracy levels needs to be combined. Our method can be readily applied to scenarios with more than two fidelity levels. It is also particularly appealing since it is complementary to other compute-efficient SBI methods. For example, Tatsuoka et al. (2025) recently proposed to train an NPE network on low-fidelity data, and to then use the resulting posterior approximation to guide sampling from the high-fidelity simulator. This approach could easily be combined with our method.
@@ Acknowledgments
The authors are grateful to Sam Power, Tim Sullivan and David Warne for helpful discussions, and to the authors of Krouglova et al. (2025) for sharing their code and identifying a bug in our implementation of their method in a preprint version of this paper. YH and AB were supported by the Research Council of Finland grant no. 362534. FXB was supported by the EPSRC grant [EP/Y022300/1]. NJ was supported by the ERC-selected UKRI Frontier Research Grant EP/Y03015X/1 and by the Simons Collaboration on Learning the Universe.
@@ A.1 Proof of Theorem˜1
Note that here, and throughout the rest of this proof, we will simplify the notation. We will drop the subscript for variances and expectations, and these should all be understood as being $u_{i}\sim\mathbb{U}$ for $i\in\{1,\ldots,m\}$ and $\theta\sim\pi$. We will also use the variable $z$ to denote a vector containing $u_{1:m}$ and $\theta$, so that $z\in\mathbb{R}^{md_{\mathcal{U}}+d_{\Theta}}$. Finally, we will write $L^{4}$ and $W^{1,4}$ to denote the spaces $L^{4}(\pi\times\mathbb{U})$ and $W^{1,4}(\pi\times\mathbb{U})$.
The proof will be structured as follows. We will express the variance of each term using the variance of individual samples. We will then use a Poincaré-type inequality to bound the variance of the objective in terms of the expected squared norm of its gradient, then we will upper bound the norm using only terms which depend on constants and terms expressing how well $G^{l-1}$ approximates $G^{l}$.
Since each term in the MLMC expansion is an MC estimator (and therefore based on independent samples), we can express the variances of each term as follows:
@@
for $l\in\{1,\ldots,L\}$. Here, we emphasise again that the expectation is over $z$, which encompasses both $\theta$ and $u_{1:m}$, and the vector $\nabla_{z}f^{l}_{\phi}(z)\in\mathbb{R}^{d_{\mathcal{U}}m+d_{\Theta}}$ for any $l\in\{1,\ldots,L\}$. This result requires the joint distribution of the prior $\pi$ and $m$ times the base measure $\mathbb{U}$ to be log-concave, which holds since the product of log-concave densities is also log-concave (since the sum of convex functions is convex) and we have assumed that the prior and base measures are independent and separately log-concave through Assumption (A1).
To simplify notation, we now introduce the vector-valued function $g^{l}(z)=g^{l}(\theta,u_{1:m})=(G^{l}_{\theta}(u_{1})^{\top},\ldots,G^{l}_{\theta}(u_{m})^{\top},\theta^{\top})^{\top}$ so that $\tilde{q}_{\phi}(x_{1:m},\theta)=\tilde{q}_{\phi}(g^{l}(\theta,u))=\tilde{q}_{\phi}(g^{l}(z))$. We will now derive our first bound, which looks at the first term in the MLMC expansion. To do so, we simplify ˜8 as follows
$\mathbb{E}\left[\left\|\nabla_{z}f_{\phi}^{0}(z)\right\|^{2}_{2}\right]$ $=\mathbb{E}\left[\left\|\nabla_{z}-\log\tilde{q}_{\phi}\left(g^{0}(z)\right)\right\|^{2}_{2}\right]$
$=\mathbb{E}\left[\left\|\nabla_{z}\log\tilde{q}_{\phi}\left(g^{0}(z)\right)\right\|^{2}_{2}\right]$
$=\mathbb{E}\left[\left\|\nabla_{z}g^{0}(z)\nabla\log\tilde{q}_{\phi}\left(g^{0}(z)\right)\right\|^{2}_{2}\right]$ (10)
$\leq\mathbb{E}\left[\left\|\nabla_{z}g^{0}(z)\right\|^{2}_{2}\left\|\nabla\log\tilde{q}_{\phi}\left(g^{0}(z)\right)\right\|^{2}_{2}\right]$ (11)
$\leq\mathbb{E}\left[\left\|\nabla_{z}g^{0}(z)\right\|^{4}_{2}\right]^{\frac{1}{2}}\mathbb{E}\left[\left\|\nabla\log\tilde{q}_{\phi}\left(g^{0}(z)\right)\right\|^{4}_{2}\right]^{\frac{1}{2}}$ (12)
Here, we have that ˜10 follows due to the chain rule, ˜11 follows due to the definition of the matrix 2-norm, ˜12 follows due to the Cauchy-Schwarz inequality of expectations.
@@
$\|\nabla_{z}g^{0}(z)\|_{2}^{2}$ $\leq\sum_{i=1}^{m}\|\nabla_{u_{i}}g^{0}(z)\|_{2}^{2}+\|\nabla_{\theta}g^{0}(z)\|^{2}_{2}$ (13)
$\leq\sum_{i=1}^{m}\sum_{j=1}^{m}\|\nabla_{u_{i}}G^{0}_{\theta}(u_{j})\|_{2}^{2}+\sum_{j=1}^{m}\|\nabla_{\theta}G^{0}_{\theta}(u_{j})\|^{2}_{2}+\|\nabla_{\theta}\theta\|_{2}^{2}$ (14)
$=\sum_{j=1}^{m}\|\nabla_{u_{j}}G^{0}_{\theta}(u_{j})\|_{2}^{2}+\sum_{j=1}^{m}\|\nabla_{\theta}G^{0}_{\theta}(u_{j})\|^{2}_{2}+d_{\Theta}$ (15)
$=\sum_{j=1}^{m}\sum_{k=1}^{d_{\mathcal{U}}}\|\nabla_{u_{jk}}G^{0}_{\theta}(u_{j})\|_{2}^{2}+\sum_{j=1}^{m}\sum_{k=1}^{d_{\Theta}}\|\nabla_{\theta_{k}}G^{0}_{\theta}(u_{j})\|^{2}_{2}+d_{\Theta}$ (16)
where ˜13 and ˜14 follow from the fact that the two norm squared of a matrix is less than the sum of the two norm squared of sub-matrices constructed through rows and columns, ˜15 follows by noticing that $\|\nabla_{\theta}\theta\|^{2}_{2}=d_{\Theta}$ and $\|\nabla_{u_{i}}G_{\theta}^{0}(u_{j})\|_{2}^{2}=0$ whenever $i\neq j$, and ˜16 follows similarly to ˜13 and ˜14. Taking squares and an expectation, we get:
$\mathbb{E}\left[\|\nabla_{z}g^{0}(z)\|_{2}^{4}\right]$
$\leq\mathbb{E}\left[\left(\sum_{j=1}^{m}\sum_{k=1}^{d_{\mathcal{U}}}\|\nabla_{u_{jk}}G^{0}_{\theta}(u_{j})\|_{2}^{2}+\sum_{j=1}^{m}\sum_{k=1}^{d_{\Theta}}\|\nabla_{\theta_{k}}G^{0}_{\theta}(u_{j})\|^{2}_{2}+d_{\Theta}\right)^{2}\right]$ (17)
$\leq(md_{\mathcal{U}}+md_{\Theta}+1)\mathbb{E}\left[\sum_{j=1}^{m}\sum_{k=1}^{d_{\mathcal{U}}}\|\nabla_{u_{jk}}G^{0}_{\theta}(u_{j})\|_{2}^{4}+\sum_{j=1}^{m}\sum_{k=1}^{d_{\Theta}}\|\nabla_{\theta_{k}}G^{0}_{\theta}(u_{j})\|^{4}_{2}+d_{\Theta}^{2}\right]$ (18)
$\leq(md_{\mathcal{U}}+md_{\Theta}+1)\left(m\|G^{0}\|^{4}_{W^{1,4}}+d_{\Theta}^{2}\right)$ (19)
@@
where ˜18 follows from $(\sum_{i=1}^{n}a_{i})^{2}\leq n\sum_{i=1}^{n}a_{i}^{2}$, ˜19 follows from the definition of the Sobolev norm and the fact that $u_{1},\ldots,u_{m}$ have the same distribution and hence the same expectation, and ˜20 follows by grouping constants together.
We now move on to bounding the second term in ˜12. Since we assumed in Assumption (A3) that $\nabla\log\tilde{q}_{\phi}$ satisfies a linear growth condition, we must have that
$\mathbb{E}\left[\|\nabla\log\tilde{q}_{\phi}(g^{0}(z))\|_{2}^{4}\right]$ $\leq\mathbb{E}\left[\left(K_{\text{grow}}(\phi)\left(\sum_{j=1}^{m}\|G_{\theta}^{0}(u_{j})\|_{2}+\|\theta\|_{2}+1\right)\right)^{4}\right]$ (21)
$\leq(m+2)^{3}K_{\text{grow}}(\phi)^{4}\mathbb{E}\left[\sum_{j=1}^{m}\|G_{\theta}^{0}(u_{j})\|^{4}_{2}+\|\theta\|_{2}^{4}+1\right]$ (22)
$\leq(m+2)^{3}K_{\text{grow}}(\phi)^{4}\left(m\|G^{0}\|_{W^{1,4}}^{4}+\mathbb{E}\left[\|\theta\|_{2}^{4}\right]+1\right)$ (23)
$\leq K_{\text{score}}(\phi)\left(\|G^{0}\|_{W^{1,4}}^{4}+1\right)$ (24)
Here, ˜21 uses the growth condition, ˜22 holds by applying $(\sum_{i=1}^{n}a_{i})^{4}\leq n^{3}\sum_{i=1}^{n}a_{i}^{4}$, ˜23 follows from the definition of Sobolev norm, and ˜24 uses the fact that $\mathbb{E}\left[\|\theta\|_{2}^{4}\right]$ is upper bounded by a constant since the expectation is against $\pi$, which a log-concave distribution and hence all its moments are finite (see Section 5.1. of Saumard and Wellner [2014]).
Combining Section˜A.1, ˜8, ˜12, ˜20, and ˜24 therefore gives: